OpenStudy (skullcandy229):

All of the following are equivalent to 2(2a + b) + 8, except _____.

OpenStudy (skullcandy229):

one of these are it 4a + 2b + 8 2(2a + b + 4) 4(a + b + 2) 4a + 2(b + 4)

OpenStudy (jango_in_dtown):

teh given expression is 2(2a+b)+8=4a+2b+8

OpenStudy (jango_in_dtown):

now check the given options

OpenStudy (jango_in_dtown):

first one is exactly the expression we got second one is 2(2a+b+4)=4a+2b+8 which is again the same expression

OpenStudy (jango_in_dtown):

check the third one its 4(a+b+2)=4a+4b+8, which is not wqual to the given expression since the coefficient of b is 4 here